# BQA: Body Language Question Answering Dataset for Video Large Language Models

### v2

@@ Abstract
A large part of human communication relies on nonverbal cues such as facial expressions, eye contact, and body language. Unlike language or sign language, such nonverbal communication lacks formal rules, requiring complex reasoning based on commonsense understanding. Enabling current Video Large Language Models (VideoLLMs) to accurately interpret body language is a crucial challenge, as human unconscious actions can easily cause the model to misinterpret their intent. To address this, we propose a dataset, BQA, a body language question answering dataset, to validate whether the model can correctly interpret emotions from short clips of body language comprising 26 emotion labels of videos of body language. We evaluated various VideoLLMs on the BQA with and without Multimodal Chain of Thought (CoT) and revealed that understanding body language is challenging, and our analyses of the wrong answers by VideoLLMs show that certain VideoLLMs made largely biased answers depending on the age group and ethnicity of the individuals. We also found the consistent error patterns in VideoLLMs. 111 The dataset is released at https://huggingface.co/datasets/naist-nlp/BQA.
@@ 1 Introduction
Our research focuses on the analysis of various emotional expressions in human body language. We created a dataset called BQA, a multiple-choice QA task, in which each body language video is associated with a question regarding a particular emotion comprising four choice answers e.g., Surprise, Confidence, Anger, and Embarrassment, reformatting Body Language Dataset created for pose estimation Luo et al. (2020). The BQA consists of 7,632 short videos (5–10 seconds, 25 fps), depicting human body language with metadata (gender, age, ethnicity) and 26 emotion labels per video. The BQA creation involves four steps using Gemini Team et al. (2024): extracting answer choices, generating questions, evaluating potential harm, and assigning difficulty labels. Moreover, we evaluated recent VideoLLMs on BQA, with and without Chain of Thought (CoT) Zhang et al. (2024b) as well as conducted multiple human evaluations, and found the task enough challenging for VideoLLMs. Error analysis revealed biases toward specific ages or ethnicities, and Multimodal CoT, despite improving accuracy, showed consistent error patterns.
@@ 2 Body Language Dataset (BoLD)
Body Language Dataset (BoLD) Luo et al. (2020) is a dataset for recognizing human actions and selecting appropriate emotions created by splitting 150 films, totaling 220 hours of footage, resulting in 9,876 video clips. Each clip is approximately 5 seconds long, comprising nearly 125 frames. These videos were annotated with 26 emotion labels Kosti et al. (2017) via crowdsourcing, where multiple annotators assigned emotion labels on a 10-point scale, which were normalized to represent the emotion of each video. Metadata, including the age, gender, and ethnicity of the individuals in the clips, is also available. However, BoLD is designed for a model directly predicting the emotion, and is not suitable for prompting with clear answers in order to investigate the capacity of VideoLLMs which expect inference with the natural language.
@@ 3 Dataset Construction
STEP1: Extract Candidates
We categorized 26 emotion labels defined in BoLD into 4 groups: Happiness, Anger, Sadness, and Pleasure, as shown in Figure 2, based on the research in which emotions are classified into four main groups James (1890). For the creation of BQA, we apply a multiple-choice question format, where the one with the highest empathy level is treated as correct, and the remaining three options are selected from different emotion groups to ensures that the choices for the QA candidates are selected from each of the groups. The example in Figure 1-1 shows that the correct answer is Suprise from the Pleasure group and the remaining candidates, i.e., Confidence, Anger and Embarrassment, are drawn from the other groups, i.e., Happiness, Anger and Sadness, respectively.
@@
Finally, as we showed Figure 1-4, we let Gemini solve the created questions by labeling each question as “Easy” when it could answer or as “Hard” if it could not answer using the prompt presented in Appendix C.2. These question labels allow us to analyze whether a hard question that is not solvable by Gemini is also difficult for other VideoLLMs. For instance, if the correct answer is “Surprise” and Gemini responds with “Anger,” then that question would be classified as “Hard.” After completing these four steps, we split the dataset into training, validation, and test sets in a 6:2:2 ratio, resulting in 4.5k, 1.5k, and 1.5k questions, respectively. The precise number of data is in Table 3 of Appendix.
@@
Our analysis of the issues labeled as "Hard" revealed the following several findings:
@@ 4 Evaluation
The models used for evaluation include VideoLLaMA2 Cheng et al. (2024), LLaVA-NeXT Zhang et al. (2024a), Qwen2-VL Wang et al. (2024), and Phi-3.5 Abdin et al. (2024) with the prompt in Appendix B.2. For the test data, we also used the proprietary models, Gemini Team et al. (2024) and GPT-4o Achiam et al. (2023). We used LoRA-Tuning Hu et al. (2022) on VideoLLaMA2 with the configuration in Appendix B.1 using the training data and let the model answer with the correct choice in a single word. All audio from the videos was removed to allow for evaluating the model’s ability to interpret body language without relying on auditory information. Additionally, we randomly selected 100 cases from the test set to measure human performance, describing the guideline in Appendix C.1. As an additional evaluation on the test set, we further employed Multimodal Chain of Thought (CoT) to make the models generate reasoning as evidence we call “rationale” for selecting answers. The prompts used for these evaluations are provided in Appendix C.2.
@@
We show the results in Table 1 with the test set and Table 4 with the valid set. GPT-4o and Gemini achieved higher accuracy than the other models. VideoLLaMA2 replied without confirming the choice format before fine-tuning (FT), resulting in a low score, but after FT, its score surpassed Gemini’s. From this result, the label assignment in STEP4 did not cause hallucinations. Regarding Gemini, which generated the questions, we found that the problems were sufficiently challenging even for the model itself. Furthermore, in STEP4 of Section 3, those labeled as Easy became unsolvable during inference, likely due to the prompt that restricted the output to single words. Regarding CoT, the results showed a large improvement, confirming its effectiveness for VideoLLMs.
@@ 6 Conclusion
Our work created a dataset called BQA to evaluate whether VideoLLMs understand body language that represent emotions and let VideoLLMs solve the questions using BQA. The results show the questions are challenging for all models, confirming the meaning of the dataset. We analyze the types of questions each model tends to get wrong, revealing that some models show a tendency to make more mistakes based on ethnicity and age group. The models also tend to focus more on the face than on body language, and we found that accuracy decreases when there are obstructions for obscuring the face. It is also essential to conduct evaluations that focus on the biases in VideoLLMs.
@@ 7.1 The Evaluation by Human Annotator
The random sampling of 100 evaluations was conducted by three people. The overall agreement score is 0.79, indicating a high level agreement. In the future, we may take care of using crowdsourcing to gather evaluations from a lot of people. Furthermore, it would be beneficial to include evaluations from people of various ethnicities to explore differing perspectives on body language that expresses emotions. However, Tedeschi et al. (2023) argues that human baselines may be unreliable due to factors such as crowdsourced worker payment issues and random sample effects. We should therefore be cautious about the baseline for our human evaluation.
@@ 8.2 License
Since BoLD does not have a clear license, we think its use for research purposes is unproblematic.
@@ A.1 Which Emotion Do VideoLLMs Often Mistake?
In Figure 6, we analyzed which labels the models predicted for the questions they answered incorrectly. The x-axis represents the correct labels, while the y-axis shows the emotions predicted by the models. The “Others” includes instances where the model output was not an emotion, such as a sentence. None of the models predicted “Pleasure” when they made mistakes, indicating a strong capability to predict actions representing “Pleasure.” However, all models frequently erred when the correct label was “Happiness,” often selecting opposing emotions like “Sadness” or “Anger.”
@@ A.2 Expansion from existing QA to Other QA Datasets
There are methods for expanding from existing QA datasets to other QA datasets. Seo et al. (2024); Nohejl et al. (2025); Ozaki et al. (2024b) created a Korean-specific NLG QA dataset by extending the existing CommonsenseQA Talmor et al. (2019); Ozaki et al. (2024a). Similarly, we extended BoLD to construct BQA, which is tailored to accommodate inputs for VideoLLMs, contributing to creating the dataset for understanding the body language of VideoLLMs.
@@ A.3 Which Gender do VideoLLMs Often Mistake?
Figure 4-A shows the tendency of each model to make mistakes based on whether the video features a male or female subject. Higher values indicate a higher likelihood of errors for the videos. From these results, we can see that none of the models exhibit bias based on gender. These findings suggest that the models are focused on human actions rather than whether the person is male or female.
@@ A.4 Trends in Generated Questions
As described in Section 3, we used Gemini, which achieved the best performance among VideoLLMs according to Fu et al. (2024), to generate the question texts. We confirmed that the generated questions predominantly begin with wh- words, i.e., What, Where, When, Who, Why, How, Which, which we think appropriate for question formation. Furthermore, we evaluated the diversity of the generated questions using Self-BLEU Zhu et al. (2018), and the results are shown in Table 2. These results are based on 4-gram analysis [^4] When calculating the score, we use nltk Bird and Loper (2004).. They indicate that the diversity of the questions is balanced across sets, confirming the fairness of the question distribution.
@@ A.7 Exhortation of Multimodal CoT
We analyzed the large accuracy improvements achieved by applying Multimodal Chain of Thought Zhang et al. (2024b). Upon examining the generated rationales, we observed that many explanations explicitly included reasons for selecting the correct answer while also providing reasons for not choosing the other options. Multimodal CoT presents the question, the answer choices, and the correct answer, then generates an explanatory rationale that derives the answer from the given choices, this rationale being subsequently used for the QA task. These rationales often clearly incorporated the correct answer within the explanation itself as we described below boxes. Unlike the traditional step-by-step reasoning approach of Chain of Thought Wei et al. (2022), this explicit inclusion of the correct answer suggests a form of answer leakage, which naturally contributes to higher accuracy. Therefore, it is important to note that the results obtained from Multimodal CoT are influenced by this answer leakage and may not fully reflect the inherent capabilities of the model.
Examples of Successful Generation about Rationale The man sitting in the armchair makes eye contact with the other man and sustains it while he speaks. This direct eye contact and attention to the conversation suggests active participation and engagement with what the other man is saying.
@@ B.1 LoRA Tuning Setting
We conducted LoRA Hu et al. (2022) Tuning with VideoLLaMA2 model. The model was trained using four NVIDIA A100-SXM4-40GB GPUs. Detailed parameters are provided in Table 5.
@@
Table 5: The hyper-parameters of VideoLLaMA2 Cheng et al. (2024) for LoRA-Tuning Hu et al. (2022) used in the experiment, and others, were set to default settings Sakai et al. (2024b). The implementation used Transformers Wolf et al. (2020).
@@ B.2 Gemini Settings
B.3 Filtering by Rule Based Algorithm
@@ B.4 The Proportion of Data Split
Figure 6: The emotional distribution of output from each model. The x-axis shows the label of the correct answer and the y-axis shows how the model got it wrong in doing so.
@@ C.1 Instruction for Human Evaluation
We conducted the human evaluation for test data of BQA. The instruction for requesting humans is shown below.
@@
C.2 The Prompt on Creating Dataset
C.3 The Content of Dataset
The structure of the metadata contained in the dataset is outlined below. The complete dataset is planned to be released at https://huggingface.co/datasets/naist-nlp/BQA.
@@ C.4 Examples of BQA
We provide several examples of actual questions and the corresponding answers from each model. The images in the Video column capture the most distinctive moments from the videos (i.e., thumbnails). The complete dataset will be available at https://huggingface.co/datasets/naist-nlp/BQA. The prompt includes only the necessary parts; for the complete details, we described in Appendix C.2.

### v3

@@ Abstract
A large part of human communication relies on nonverbal cues such as facial expressions, eye contact, and body language. Unlike language or sign language, such nonverbal communication lacks formal rules, requiring complex reasoning based on commonsense understanding. Enabling current Video Large Language Models (VideoLLMs) to accurately interpret body language is a crucial challenge, as human unconscious actions can easily cause the model to misinterpret their intent. To address this, we propose a dataset, BQA, a body language question answering dataset, to validate whether the model can correctly interpret emotions from short clips of body language comprising 26 emotion labels of videos of body language. We evaluated various VideoLLMs on the BQA with and without Multimodal Chain of Thought (CoT) and revealed that understanding body language is challenging, and our analyses of the wrong answers by VideoLLMs show that certain VideoLLMs made largely biased answers depending on the age group and ethnicity of the individuals. We also found consistent error patterns in VideoLLMs111 The dataset is available at https://huggingface.co/datasets/naist-nlp/BQA. .
@@ 1 Introduction
Our research focuses on the analysis of various emotional expressions in human body language. We created a dataset called BQA, a multiple-choice QA task, in which each body language video is associated with a question regarding a particular emotion comprising four choice answers, e.g., Surprise, Confidence, Anger, and Embarrassment, reformatting the Body Language Dataset created for pose estimation Luo et al. (2020). The BQA consists of 7,632 short videos (5–10 seconds, 25 fps), depicting human body language with metadata (gender, age, ethnicity) and 26 emotion labels per video. The BQA creation involves four steps using Gemini Team et al. (2024): extracting answer choices, generating questions, evaluating potential harm, and assigning difficulty labels. Moreover, we evaluated recent VideoLLMs on BQA, with and without Chain of Thought (CoT) Zhang et al. (2024b) as well as conducted multiple human evaluations, and found the task enough challenging for VideoLLMs. Error analysis revealed biases toward specific ages or ethnicities, and Multimodal CoT, despite improving accuracy, showed consistent error patterns.
@@ 2 Body Language Dataset (BoLD)
Body Language Dataset (BoLD) Luo et al. (2020) is a dataset for recognizing human actions and selecting appropriate emotions created by splitting 150 films, totaling 220 hours of footage, resulting in 9,876 video clips. Each clip is approximately 5 seconds long, comprising nearly 125 frames. These videos were annotated with 26 emotion labels Kosti et al. (2017) via crowdsourcing, where multiple annotators assigned emotion labels on a 10-point scale, which were normalized to represent the emotion of each video. Metadata, including the age, gender, and ethnicity of the individuals in the clips, is also available. However, BoLD is designed for a model directly predicting the emotion, and is not suitable for prompting with clear answers in order to investigate the capacity of VideoLLMs, which expect inference with natural language.
@@ 3 Dataset Construction
STEP1: Extract Candidates
We categorized 26 emotion labels defined in BoLD into 4 groups: Happiness, Anger, Sadness, and Pleasure, as shown in Figure 2, based on the research in which emotions are classified into four main groups James (1890). For the creation of BQA, we apply a multiple-choice question format, where the one with the highest empathy level is treated as correct, and the remaining three options are selected from different emotion groups to ensure that the choices for the QA candidates are selected from each of the groups. The example in Figure 1-1 shows that the correct answer is Surprise from the Pleasure group and the remaining candidates, i.e., Confidence, Anger, and Embarrassment, are drawn from the other groups, i.e., Happiness, Anger, and Sadness, respectively.
@@
Finally, as shown in Figure 1-4, we let Gemini solve the created questions by labeling each question as “Easy” when it could answer or as “Hard” if it could not answer using the prompt presented in Appendix C.3. These question labels allow us to analyze whether a hard question that is not solvable by Gemini is also difficult for other VideoLLMs. For instance, if the correct answer is “Surprise” and Gemini responds with “Anger,” then that question would be classified as “Hard.” After completing these four steps, we split the dataset into training, validation, and test sets in a 6:2:2 ratio, resulting in 4.5k, 1.5k, and 1.5k questions, respectively. The precise number of data is in Table 3.
@@
Our analysis of the issues labeled as "Hard" revealed the following findings:
@@ 4 Evaluation
The models used for evaluation include VideoLLaMA2 Cheng et al. (2024), LLaVA-NeXT Zhang et al. (2024a), Qwen2-VL Wang et al. (2024), and Phi-3.5 Abdin et al. (2024) with the prompt in Appendix B.2. For the test data, we also used the proprietary models, Gemini Team et al. (2024) and GPT-4o OpenAI et al. (2024). We used LoRA-Tuning Hu et al. (2022) on VideoLLaMA2 with the configuration in Appendix B.1 using the training data and let the model answer with the correct choice in a single word. All audio from the videos was removed to allow for evaluating the model’s ability to interpret body language without relying on auditory information. Additionally, we randomly selected 100 cases from the test set to measure human performance, describing the guideline in Appendix C.1. As an additional evaluation on the test set, we further employed Multimodal Chain of Thought (CoT) to make the models generate reasoning as evidence we call “rationale” for selecting answers. The prompts used for these evaluations are provided in Appendix C.3.
@@
We show the results in Table 1 with the test set and Table 4 with the valid set. GPT-4o and Gemini achieved higher accuracy than the other models. VideoLLaMA2 replied without confirming the choice of format before fine-tuning (FT), resulting in a low score, but after FT, its score surpassed Gemini’s. From this result, the label assignment in STEP4 did not cause hallucinations. Regarding Gemini, which generated the questions, we found that the problems were sufficiently challenging even for the model itself. Furthermore, in STEP4 of Section 3, those labeled as Easy became unsolvable during inference, likely due to the prompt that restricted the output to single words. Regarding CoT, the results showed a large improvement, confirming its effectiveness for VideoLLMs.
@@ 6 Conclusion
Our work created a dataset called BQA to evaluate whether VideoLLMs understand body language that represents emotions and let VideoLLMs solve the questions using BQA. The results show the questions are challenging for all models, confirming the meaning of the dataset. We analyze the types of questions each model tends to get wrong, revealing that some models show a tendency to make more mistakes based on ethnicity and age group. The models also tend to focus more on the face than on body language, and we found that accuracy decreases when there are obstructions that obscure the face. It is also essential to conduct evaluations that focus on the biases in VideoLLMs.
@@ 7.1 The Evaluation by Human Annotator
The random sampling of 100 evaluations was conducted by three people. The overall agreement score is 0.79, indicating a high level of agreement. In the future, we may take care of using crowdsourcing to gather evaluations from a lot of people. Furthermore, it would be beneficial to include evaluations from people of various ethnicities to explore differing perspectives on body language that expresses emotions. However, Tedeschi et al. (2023) argues that human baselines may be unreliable due to factors such as crowdsourced worker payment issues and random sample effects. We should therefore be cautious about the baseline for our human evaluation.
@@ 8.2 License
Since BoLD does not have a clear license, we believe its use for research purposes is unproblematic.
@@ A.1 Which Emotion Do VideoLLMs Often Mistake?
Figure 5: The emotional distribution of output from each model. The x-axis shows the label of the correct answer, and the y-axis shows how the model got it wrong in doing so.
We analyzed which labels the models predicted for the questions they answered incorrectly in Figure 5. The x-axis represents the correct labels, while the y-axis shows the emotions predicted by the models. The “Others” includes instances where the model output was not an emotion, such as a sentence. None of the models predicted “Pleasure” when they made mistakes, indicating a strong capability to predict actions representing “Pleasure.” However, all models frequently failed when the correct label was “Happiness,” often selecting opposing emotions like “Sadness” or “Anger.”
@@ A.2 Expansion from existing QA to Other QA Datasets
There are several methods for extending existing QA datasets to new tasks or modalities. For example, Seo et al. (2024) created KoCommonGEN v2 as an expansion of Korean CommonGen Seo et al. (2022), and Lynn et al. (2025) constructed a new culturally-aware dataset based on Belebele Bandarkar et al. (2024). Demszky et al. (2018) transformed existing QA datasets, such as SQuAD Rajpurkar et al. (2016) and MovieQA Tapaswi et al. (2016), into an NLI dataset, highlighting how adapting existing resources can yield valuable contributions. Similarly, Sakajo et al. (2025) developed Tonguescape to benchmark VideoLLMs. Building on this trend, we extend BoLD to construct BQA, a dataset designed to serve as input for VideoLLMs and support the understanding of body language.
@@ A.3 Which Gender do VideoLLMs Often Mistake?
Figure 4-A shows the tendency of each model to make mistakes based on whether the video features a male or female subject. Higher values indicate a higher likelihood of errors for the videos. From these results, we can see that none of the models exhibits bias based on gender. These findings suggest that the models are focused on human actions rather than whether the person is male or female.
@@ A.4 Trends in Generated Questions
As described in Section 3, we used Gemini, which achieved the best performance among VideoLLMs according to Fu et al. (2024), to generate the question texts. We confirmed that the generated questions predominantly begin with wh-words, i.e., What, Where, When, Who, Why, How, Which, which we believe is appropriate for question formation. Furthermore, we evaluated the diversity of the generated questions using Self-BLEU Zhu et al. (2018), and the results are shown in Table 2. These results are based on 4-gram analysis [^4] When calculating the score, we use nltk Bird and Loper (2004).. They indicate that the diversity of the questions is balanced across sets, confirming the fairness of the question distribution.
@@ A.7 Exhortation of Multimodal CoT
We analyzed the large accuracy improvements achieved by applying Multimodal Chain of Thought Zhang et al. (2024b). Upon examining the generated rationales, we observed that many explanations explicitly included reasons for selecting the correct answer while also providing reasons for not choosing the other options. Multimodal CoT presents the question, the answer choices, and the correct answer, then generates an explanatory rationale that derives the answer from the given choices, this rationale being subsequently used for the QA task. These rationales often clearly incorporated the correct answer within the explanation itself, as we described below the boxes. Unlike the traditional step-by-step reasoning approach of Chain of Thought Wei et al. (2022), this explicit inclusion of the correct answer suggests a form of answer leakage, which naturally contributes to higher accuracy. Therefore, it is important to note that this answer leakage influences the results obtained from Multimodal CoT and may not fully reflect the inherent capabilities of the model.
Examples of Successful Generation about Rationale The man sitting in the armchair makes eye contact with the other man and sustains it while he speaks. This direct eye contact and attention to the conversation suggest active participation and engagement with what the other man is saying.
@@ B.1 LoRA Tuning Setting
We conducted LoRA Hu et al. (2022) tuning with the VideoLLaMA2 model. The model was trained using four NVIDIA A100-SXM4-40GB GPUs. Detailed parameters are provided in Table 5.
@@
Table 5: The hyperparameters of VideoLLaMA2 Cheng et al. (2024) for LoRA-Tuning Hu et al. (2022) used in the experiment, and others, were set to default settings Sakai et al. (2024b). The implementation used the Transformers library Wolf et al. (2020).
@@ B.2 Gemini Settings
B.3 Filtering by Rule-Based Algorithm
@@ C.1 Instruction for Human Evaluation
We conducted human evaluations of the BQA test data. The instruction for requesting humans is shown below.
@@
C.2 The Content of the Dataset
The structure of the metadata contained in the dataset is outlined below. The complete dataset is released at https://huggingface.co/datasets/naist-nlp/BQA.
@@ C.4 Examples of BQA
We provide several examples of actual questions and the corresponding answers from each model. The images in the Video column capture the most distinctive moments from the videos (i.e., thumbnails). The complete dataset is available at https://huggingface.co/datasets/naist-nlp/BQA. The prompt includes only the necessary parts; for the complete details, we describe in Appendix C.3.